OpenStudy (anonymous):

prove if u can that 0! is equal to 1?

OpenStudy (anonymous):

0! is defined as 1. QED.

OpenStudy (anonymous):

prove it

OpenStudy (anonymous):

I think you misunderstand the terms ' definition ' and 'prove'.

OpenStudy (anonymous):

You can explain why it 'makes sense' it should be one, but when it comes down it it, it's just defined to be that.

OpenStudy (anonymous):

everything has a prove in mathematics ,

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Axiom . Now shut up please.

OpenStudy (anonymous):

see, i can prove it but there is always some "bug". maybe it will work for you but for some other person (who has more knowledge of maths) can prove it wrong.

OpenStudy (anonymous):

Prove it.

OpenStudy (anonymous):

I mean, like I said there are EXPLANATIONS like this http://www.adonald.btinternet.co.uk/Factor/Zero.html But it is different.

OpenStudy (anonymous):

prove it ,it does,nt mater ,it will add knowlede

OpenStudy (anonymous):

its a convention

OpenStudy (anonymous):

lots of math people got together and decided, hey this is going to be the way

OpenStudy (anonymous):

but the reason for it , it simplifies many formulas

OpenStudy (anonymous):

thanks ,newton it was helpful

OpenStudy (anonymous):

so maybe you should ask, whats the reason for the convention, thats a better question,

OpenStudy (anonymous):

it can be proved using Gamma function

OpenStudy (anonymous):

its just a definition, so you cant ask to prove a definition

OpenStudy (anonymous):

or simply using the definition of the n!

OpenStudy (anonymous):

you cant prove for example, that a bachelor is unmarried. its a definition

OpenStudy (anonymous):

"It can be proved using the definition" Hmm, sounds like a proof to me (xD)

OpenStudy (anonymous):

well n! = n * n-1 *...*3*2*1

OpenStudy (anonymous):

theres isnt 0! in it

OpenStudy (anonymous):

hehe

OpenStudy (anonymous):

i am on my phone so its hard for me to type the whole derivation. or i would have given you

OpenStudy (anonymous):

proof, by definition? ok i guess

OpenStudy (anonymous):

but we dont define things just willy nilly. they are usually well good god reasons for it

OpenStudy (anonymous):

n!=n(n-1! put n=1

OpenStudy (anonymous):

While you're at it, prove all the axioms please, umza? And (technically), \[1! = 1 \times 0! \] \[2! = 2 \times 1! \] \[3! = 3 \times 2! \] ... \[n! = n \times (n-1)! \]

OpenStudy (anonymous):

Umza please go learn the definition of 'proof' and come back.

OpenStudy (anonymous):

mean?INew

OpenStudy (anonymous):

uzma, that doesnt work

OpenStudy (anonymous):

ok thank u

OpenStudy (anonymous):

uzma , what about 0! = 0 * (0-1)! ?

OpenStudy (anonymous):

anyways, i like the binomial series , i think its called pochhammer

OpenStudy (anonymous):

(0-1)! does it work?

OpenStudy (anonymous):

the reason why we have 0! is because we have n choose r = n! / ( n-r)! r!

OpenStudy (anonymous):

and when you have r = 0, we have n! / n! 0!

OpenStudy (anonymous):

the actual proof is done using Gamma function

OpenStudy (anonymous):

wait, its false that 0! = 0 * (0-1)!

OpenStudy (anonymous):

the proof for what, the bionomial series

OpenStudy (anonymous):

oh sorry, i misread the question and thought that you are asking to prove 0 = 1 instead of 0! = 1 lol. its very easy to prove.

OpenStudy (anonymous):

OMG yes it's SOOOO EASY to prove - just do it quickly please?

OpenStudy (anonymous):

but i don't know to type, its not easy to type "please".

OpenStudy (anonymous):

Oh, I thought so.

OpenStudy (anonymous):

i'm back!!! are you guys there. lol